# UniAttn: Reducing Inference Costs via Softmax Unification for Post-Training LLMs

### v1

@@ 5 Conclusion
In this work, we explored LLM redundancies and efficient architectures in the context of post-training to reduce the inference costs of LLMs. Theoretically, we identified that the existing cross-layer KV sharing architecture diminishes the depth factor in pre-trained LLMs, potentially limiting their ability to acquire new capabilities during post-training. To address this limitation, we investigated the primary bottleneck in LLMs—the Softmax operation—and discovered significant redundancies across various pre-trained LLMs and datasets. Based on these findings, we proposed UniAttn, an efficient architecture that leverages these redundancies while preserving LLM capabilities through linear compensation in post-training. Extensive experiments on diverse pre-trained models and post-training datasets validate the effectiveness of UniAttn. Our UniAttn is particularly well-suited for real-world deployments, enjoying faster inference and less memory cost.
Appendix A More Implementation Details for Reproducibility
Post-Training Datasets. Regarding the medical domain, we directly employ the instruction tuning dataset for PMC [^1] https://huggingface.co/datasets/axiong/pmc_llama_instructions (Wu et al., 2023) (i.e., training stage-2 for PMC-LLaMA) for post-training. The dataset consists of 513,999 instruction-input-output pairs. Regarding the general instruction tuning scenario, we employ the Tulu 3 SFT Mixture [^2] https://huggingface.co/datasets/allenai/tulu-3-sft-mixture and filter for data with only 1 round of conversation, resulting in 896,070 input-output pairs.
Training Hyperparameters. The training hyperparameters for all experiments are reported in Table 4, Table 5, and Table 6. Note that we adopt the learning rate and weight decay values according to existing research, in which those pre-trained models are post-trained under reported schedules.
@@ Appendix A More Implementation Details for Reproducibility
    "Below is an instruction that describes a task, paired with an input that
    provides further context. Write a response that appropriately completes the
@@
Appendix B More Experiment Results
@@
Table 9: Performance (%) comparison with similarity-based Superblock grouping methods.
Method Similarity Based Grouping # of Operating Layers PubMedQA MedMCQA MedQA AVG
UniAttn $\times$ 3 78.2 59.2 62.1 66.5
$\surd$ 3 77.6 57.3 57.8 64.2
UniAttn $\times$ 6 77.2 58.1 62.9 66.1
$\surd$ 6 72.0 55.6 58.8 62.1
UniAttn $\times$ 9 77.4 56.5 60.4 64.8
$\surd$ 9 77.4 54.2 56.0 62.5
UniAttn $\times$ 12 79.0 57.6 59.5 65.4
$\surd$ 12 77.0 55.9 54.8 62.6
Different Compensation Designs. To compare different compensation designs, we attach the input and output of the $W_{c}$ transformation to different activations. As shown in Table 8, directly compensating the MHA output with its input yields the best results. When adopting “finer compensation granularities”, i.e., positioning the input and output closer together (Rows 3 and 4), post-training performance decreases. Similarly, adopting “coarser compensation granularities”, i.e., moving the output compensation to after the FFN (Row 2), also results in lower performance. These findings suggest that using a linear transformation to compensate for errors specifically within the MHA module is the optimal approach. This aligns with our analysis in Section 3.3, which shows that the compensation error within the MHA module is significantly smaller in magnitude than its expectation.
Similarity-based Superblock Grouping. We employed a simple fixed grouping method for Superblock construction, grouping every 4 layers in the top half layers as a Superblock. To investigate the impact of different Superblock grouping strategies, we experimented with a similarity-based grouping scheme. In this scheme, layers with the most similar Softmax activations, as measured by cosine similarity (Figure 2), are grouped together, and a linear compensation matrix is added to layers utilizing activations from other layers. As shown in Table 9, the similarity-based grouping consistently underperforms the fixed-size Superblock grouping (default implementation) across all settings. This suggests that the simple fixed-size Superblock approach is sufficient to achieve strong post-training performance.
Appendix C Pipeline for Applying UniAttn during Post-Training
We provide a detailed pipeline for applying UniAttn as Algorithm 1.
@@ Appendix C Pipeline for Applying UniAttn during Post-Training
        Calculate the optimal initialization for the $W_{c}$ matrix in layer $i_{s}+(i-1)b+j$ according to Theorem 3.7.
        Insert the initialized $W_{c}$ matrix to layer $i_{s}+(i-1)b+j$ in model $M_{\text{uni}}$.
@@
  Freeze all training parameters other than the $W_{c}$ matrices in $M_{\text{uni}}$. Train $W_{c}$ with dataset $D$ and apply early stop when loss stops to decrease (we consider the exponential moving average of training loss).
  Conduct full fine-tune on $M_{\text{uni}}$ with $D$.
@@
Appendix D Proof for Lemma 3.1
@@
As $||\text{Norm}(\mathbf{x}_{k})||=1$ for all $k$, then:
$||\mathbf{x}_{L}||$ $=||\mathbf{x}_{0}+\sum_{k=1}^{L}F_{k}(\text{Norm}(\mathbf{x}_{k-1}))||$ (18)
$\leq||\mathbf{x}_{0}||+\sum_{k=1}^{L}||F_{k}(\text{Norm}(\mathbf{x}_{k-1}))||$
@@
$=||\mathbf{x}_{0}||+\lambda L$
Appendix E Insights for 3.5
With Lemma 3.1 and Proposition 3.2, we have shown that pre-trained Pre-Norm LLMs can be approximated as linear Pre-Norm systems. Suppose a pre-trained LLM operates on activations $\mathbf{x}\in\mathbb{R}^{d}$, each layer $i$ in the pre-trained LLM can be treated as a linear transformation matrix $W_{i}\in\mathbb{R}^{d\times d}$. A generic gradient-based update can be expressed as:
@@
$||W_{i}^{t+1}-W_{i}^{0}||=\sqrt{tr[(||W_{i}^{t+1}-W_{i}^{0}||)(||W_{i}^{t+1}-W_{i}^{0}||)^{T}]}=\sqrt{tr(U\Sigma V^{T}V\Sigma^{T}U^{T})}=\sqrt{tr(\Sigma\Sigma^{T})}=\sqrt{\sum_{i=1}^{d}\sigma_{i}^{2}}$ (21)
@@
$||W_{i}^{t+1}-W_{i}^{0}||=\sqrt{\sum_{i=1}^{d}\sigma_{i}^{2}}\approx\sigma_{\max}(W_{i}^{t+1}-W_{i}^{0})$ (22)
@@
$\sigma_{\max}(W_{i}^{t+1})\leq\sigma_{\max}(W_{i}^{0})+\sigma_{\max}(W_{i}^{t+1}-W_{i}^{0})\leq\sum_{t=0}^{t}||\eta\nabla_{W_{i}}\mathcal{L}^{t}||+\sigma_{\max}(W_{i}^{0})$ (23)
Due to having smaller values of $||\eta\nabla_{W_{i}}\mathcal{L}^{t}||$ (i.e., smaller gradient norm), compared to earlier layers, the top layers typically have smaller updates on the largest singular value, thus tend to have smaller $\sigma_{\max}$ values. This further leads to 3.5 we propose in the article.
@@
Solving for the optimal initialization of $W_{c}$ is equivalent to solving the following optimization problem:
$\min_{W_{c}}\ ||\mathbb{E}(\mathbf{x}_{i+b}W_{c}-\epsilon)||^{2}$ (24)
Since $W_{c}$ is the fixed weight of the linear transformation applied to input, $W_{c}$ is independent of $\mathbf{x}_{i+b}$. Therefore, we can re-write the optimization problem as:
$\min_{W_{c}}\ ||\mathbb{E}(\mathbf{x}_{i+b})W_{c}-\mathbb{E}(\epsilon)||^{2}$ (25)
The solution to the above system is given by $W_{c}=\mathbb{E}(\mathbf{x}_{i+b})^{+}\mathbb{E}(\epsilon)$, where $\mathbb{E}(\mathbf{x}_{i+b})^{+}$ denotes the Moore–Penrose inverse of $\mathbb{E}(\mathbf{x}_{i+b})$. By plugging in the closed-form expression of $\mathbb{E}(\mathbf{x}_{i+b})^{+}$, we achieve:
@@
where $\mathbb{E}(\mathbf{x}_{i+b})=U\Sigma V^{T}$ denotes the SVD decomposition of $\mathbb{E}(\mathbf{x}_{i+b})$. ∎
Appendix G More Insights on Linear Compensation

### v2

@@ 5 Conclusion
In this work, we explored LLM redundancies and efficient architectures in the context of post-training to reduce the inference costs of LLMs. While existing KV-cache sharing methods successfully reduce the memory overhead, they yield limited inference time reduction during deployment. To address this limitation, we investigated the primary bottleneck in LLMs, the Softmax operation, and discovered significant redundancies across various pre-trained LLMs and datasets. Based on these findings, we proposed UniAttn, an efficient architecture that leverages these redundancies while preserving LLM capabilities through linear compensation in post-training. We further support UniAttn with a theoretical analysis, demonstrating its advantage over baselines. Extensive experiments on a series of models and datasets validate the effectiveness of UniAttn. Our UniAttn is particularly well-suited for real-world deployments, enjoying faster inference and less memory cost.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
Appendix A License for Scientific Artifacts
The Training dataset PMC (Wu et al., 2023) is licensed under OpenRAIL license [^1] https://www.licenses.ai/blog/2022/8/26/bigscience-open-rail-m-license. The Training dataset TuLu3 (Lambert et al., 2024) is licensed under Open Data Commons License [^2] https://opendatacommons.org/licenses/. The Mistral model (Jiang et al., 2023), Gemma-2 model (Rivière et al., 2024), LLaMA-2 (Huang et al., 2023) model, and LLaMA-3.1 model(Dubey et al., 2024) are licensed under Apache License 2.0333https://choosealicense.com/licenses/apache-2.0/. The LongAlign evaluation dataset is also licensed under Apache License 2.0444https://choosealicense.com/licenses/apache-2.0/. The evaluation datasets (Hsieh et al., 2024; Talmor et al., 2019; Sap et al., 2019; Jin et al., 2020; Pal et al., 2022; Jin et al., 2019) are subject to the MIT license [^5] https://choosealicense.com/licenses/mit/. All usages of scientific artifacts in this paper obey the corresponding licenses.
Appendix B More Implementation Details for Reproducibility
Post-Training Datasets. Regarding the medical domain, we directly employ the instruction tuning dataset for PMC [^6] https://huggingface.co/datasets/axiong/pmc_llama
_instructions
(Wu et al., 2023) (i.e., training stage-2 for PMC-LLaMA) for post-training. The dataset consists of 513,999 instruction-input-output pairs. Regarding the general instruction tuning scenario, we employ the Tulu 3 SFT Mixture [^7] https://huggingface.co/datasets/allenai/tulu-3-sft-mixture and filter for data with only 1 round of conversation, resulting in 896,070 input-output pairs.
Training Hyperparameters. The training hyperparameters for all experiments are reported in table Table 6, Table 7 and Table 8. Note that we adopt the learning rate and weight decay values according to existing research, in which those pre-trained models are post-trained under reported schedules.
@@ Appendix B More Implementation Details for Reproducibility
    "Below is an instruction that
    describes a task, paired with an input
    that provides further context. Write a
    response that appropriately completes
[…]
@@
Experiment Packages. We report the version numbers of used packages in Table 9.
Appendix C More Experiment Results
Quantitative Analysis of Softmax Redundancies. In Figure 2 of the main article, we observed that Softmax activations in top half layers share a high cross-layer similarity across different models and datasets. To provide a more comprehensive analysis, here we take LLaMA-3.1 8B as an example and present some statistics of the cosine similarity of Softmax activations. As shown in Table 11, cosine similarities appear to be consistent across different datasets, with average values ranging from 0.79 to 0.83. The maximum cosine similarity remains consistently high at 0.97, indicating that the overall range of cosine similarity remains stable across layers and datasets. The differences for cosine similarity between the top and bottom halves of layers indicate that Softmax operations are more redundant in deeper layers, as noted in the main article. The standard deviation values suggest that redundancy is fairly consistent across datasets. Overall, these results indicate that Softmax activations for pre-trained LLMs exhibit high redundancies.
@@ Appendix C More Experiment Results
Dataset Average Difference Between Top Half and Bottom Half Layers Max Min Standard Deviation
PMC 0.8347 0.1314 0.9729 0.0159 0.1955
Math 0.8243 0.1149 0.9734 0.0384 0.1860
Tulu3 0.8214 0.1212 0.9719 0.0307 0.1896
F15K 0.7927 0.1475 0.9710 0.0033 0.2109
Table 11: Statistics of the cosine similarity of Softmax activations for LLaMA-3.1 8B across different datasets.
@@
[…]
Different Compensation Designs. To compare different compensation designs, we attach the input and output of the $W_{c}$ transformation to different activations. As shown in Table 12, directly compensating the MHA output with its input yields the best results. When adopting “finer compensation granularities”, i.e., positioning the input and output closer together (Rows 3 and 4), post-training performance decreases. Similarly, adopting “coarser compensation granularities”, i.e., moving the output compensation to after the FFN (Row 2), also results in lower performance. These findings suggest that using a linear transformation to compensate for errors specifically within the MHA module is the best-performing approach.
@@
   Calculate the initialization for the $W_{c,i_{s}+(i-1)b+j}$ matrix in layer $i_{s}+(i-1)b+j$ according to Theorem 3.1.
   Insert the initialized $W_{c,i_{s}+(i-1)b+j}$ matrix to layer $i_{s}+(i-1)b+j$ in model $M_{\text{uni}}$.
@@
 Freeze all training parameters other than the $W_{c}$ matrices in $M_{\text{uni}}$. Train $W_{c}$ with dataset $D$ and apply early stop when loss stops decreasing (we consider the exponential moving average of training loss).
 Conduct full fine-tuning on $M_{\text{uni}}$ with $D$.
@@
Appendix D Pipeline for Applying UniAttn during Post-Training
We provide a detailed pipeline for applying UniAttn as Algorithm 1.
Appendix E Theoretical Insights for Proposing 3.2
We first give a series of conclusions (namely Lemma E.1, Proposition E.2, Proposition E.3, and Proposition E.4) as preliminaries.
Lemma E.1.
(A linear Pre-Norm system has bounded growth) $f_{i}:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$, $i\in\mathbb{N}$ are LINEAR transformations and $\text{Norm}(\cdot)$ yields a vector with unit Frobenius-norm (denoted as $||\cdot||$). Let $\mathbf{x}_{0}\in\mathbb{R}^{n}$ be the input to the system. Consider an $L$-layer Pre-Norm architecture defined by:
$\mathbf{x}_{k}=\mathbf{x}_{k-1}+f_{k}(\text{Norm}(\mathbf{x}_{k-1})),$ (10)
$\quad\text{for}\quad k=1,2,\dots,L.$
[…]
@@ Appendix D Pipeline for Applying UniAttn during Post-Training
As $\text{Norm}(\mathbf{x}_{k})||=1$ for all $k$, then:
[…]
@@
[…]
Lemma E.1 shows that a Pre-Norm linear system has bounded growth. The norm of output by each layer grows at most linearly with depth $L$.
Proposition E.2.
(Pre-trained LLMs are linear Pre-Norm systems) Pre-trained decoder-based LLMs exhibit a high linearity score. Formally, let $A,B\in\mathbb{R}^{n\times d}$ denote the normalized input and output of a decoder block in LLM, respectively,
[…]
Proposition E.2 has been validated by (Razzhigaev et al., 2024) on pre-trained LLMs, suggesting that we can approximate each decoder block by a linear transformation with bounded singular values, thus placing pre-trained LLMs structurally “close” to a linear Pre-Norm system as in Lemma E.1. In the following discussion, we approximate pre-trained LLMs with an equivalent linear Pre-Norm system and investigate its input and output in each layer.
Proposition E.3.
(Pre-trained LLM layers primarily change the magnitude of activations) The input and output of a layer in pre-trained decoder-based LLMs share a high cosine similarity.
Proposition E.3 has been validated by (Men et al., 2024) on pre-trained LLMs, indicating that each layer in LLMs predominantly change the magnitude of the activation. We leverage the following conclusion for analyzing the change of activation magnitude.
Proposition E.4.
(Top layers of Pre-Norm transformers have smaller singular values) Training Pre-Norm transformers leads to top layers receiving smaller updates (gradient norms decrease from bottom to top).
(Xiong et al., 2020) has given both a theoretical proof sketch and experimental evidence of Proposition E.4. Based on Proposition E.2, we approximate each layer in LLMs as a linear transformation. To further show that the approximated linear transformations of the top layers in Pre-Norm LLMs tend to have smaller singular values, we give the following proof:
Proof.
With Lemma E.1 and Proposition E.2, we have shown that pre-trained Pre-Norm LLMs can be approximated as linear Pre-Norm systems. Suppose a pre-trained LLM operates on activations $\mathbf{x}\in\mathbb{R}^{d}$, each layer $i$ in the pre-trained LLM can be treated as a linear transformation matrix $W_{i}\in\mathbb{R}^{d\times d}$. A generic gradient-based update can be expressed as:
@@
$||$ $W_{i}^{t+1}-W_{i}^{0}||$ (17)
[…]
$=\sqrt{tr(U\Sigma V^{T}V\Sigma^{T}U^{T})}$
$=\sqrt{tr(\Sigma\Sigma^{T})}=\sqrt{\sum_{i=1}^{d}\sigma_{i}^{2}}$
@@
$W_{i}^{t+1}-W_{i}^{0}||=\sqrt{\sum_{i=1}^{d}\sigma_{i}^{2}}\approx\sigma_{\max}(W_{i}^{t+1}-W_{i}^{0})$ (18)
@@
$\sigma_{\max}(W_{i}^{t+1})$ $\leq\sigma_{\max}(W_{i}^{0})+\sigma_{\max}(W_{i}^{t+1}-W_{i}^{0})$ (19)
[…]
Due to having smaller values of $\eta\nabla_{W_{i}}\mathcal{L}^{t}||$ (i.e., smaller gradient norm), compared to earlier layers, the top layers typically have smaller updates on the largest singular value, thus tend to have smaller $\sigma_{\max}$ values. ∎
Since pre-trained LLMs mainly change the magnitude of the activation (as shown in Proposition E.3), we can use the size of the largest singular values in each layer as an indicator of the corresponding activation magnitude change. In Pre-Norm architectures, each layer operates on an input with a constant norm. This leads to the assumption that the top-layer transformations generally generate outputs with smaller norms, which is 3.2 we propose in the article.
@@
Solving for the initialization of $W_{c}$ that incurs minimal compensation error is equivalent to solving the following optimization problem:
[…]
Since we use a small subset of the training dataset (with $S$ samples, we denote this as the initialization set) to calculate the initialization, the objective above can be rewritten as:
$\min_{W_{c}}\ \frac{1}{S}\sum_{j=1}^{S}||\mathbf{x}_{i+b,j}W_{c}-\epsilon_{j}||||.$ (21)
This is a Least Absolute Deviations problem, and can be solved with the simplex algorithm, which is NP as shown in (Fearnley and Savani, 2015). To guarantee that the initialization method is efficient even in the worst case, we in turn consider the Least Squares variant:
[…]
The Least Squares variant is a convex optimization problem, and has a closed-form solution as follows:
$W_{c}=X^{+}E,$ (23)
where $X$ is a matrix created by concatenating all $\mathbf{x}_{i+b,j}$ instances, $E$ is created by concatenating all $\epsilon_{j}$ instances, and $X^{+}$ represents the Moore-Penrose inverse of $X$. When actually caculating for $W_{c}$, it is time-consuming to calculate $X^{+}$ as the size of $X$ is big. Therefore, we set the maximum instance size of $X$ as $v$. Specifically, we average $\mathbf{x}$’s and $\epsilon$’s for every $\frac{S}{v}$ instances in the initialization set, and calculate $W_{c}$ on the averaged $\mathbf{x}$’s and $\epsilon$’s. For simplicity, we denote the averaged $\mathbf{x}$’s and $\epsilon$’s as $\mathbb{E}(\mathbf{x}_{i+b})$ and $\mathbb{E}(\epsilon)$. Consequently, the calculation for $W_{c}$ is formulated as:
@@
where $V\Sigma^{+}U^{T}$ is the Moore-Penrose inverse of $\mathbb{E}(\mathbf{x}_{i+b})$, calculated via SVD. ∎
$v$ N/A 1 2 3 4 8 16
PPL 5.914 2.890 2.871 2.835 2.850 2.822 2.837
Table 13: Perplexity results for initializations with different $v$ values on LLaMA-3.1 8B in the medical domain. “N/A” represents applying Softmax Unification without linear projections.
To find an appropriate $v$ value for initialization, we have conducted preliminary experiments on LLaMA-3.1 8B in the medical domain. Specifically, we have benchmarked the perplexity on the initialization set after inserting all linear projections with initialized weights calculated. As shown in Table 13, most performance contribution comes from the act of adding initialization, and increasing the number of $v$ affects the PPL values slightly. These experiments guided us to adopt $v=1$ for our experiments due to its simplicity.
Appendix G Proof for Proposition 3.3
Proof.
$\mathbf{x}^{\prime}_{i+1}$ $=\text{softmax}(\frac{\mathbf{x}_{i+1}W_{q,i+1}K^{T}_{i}}{\sqrt{d_{k}}})V_{i}W_{o,i+1}$ (25)
$\quad+\mathbf{x}_{i+1}$
$=\text{softmax}(\frac{(\mathbf{x}_{i}+\delta_{i})W_{q,i+1}K^{T}_{i}}{\sqrt{d_{k}}})V_{i}W_{o,i+1}$
$\quad+\mathbf{x}_{i+1}$
$=\text{softmax}(\frac{\mathbf{x}_{i}W_{q,i+1}K^{T}_{i}}{\sqrt{d_{k}}}$
$\quad+\frac{\delta_{i}W_{q,i+1}K^{T}_{i}}{\sqrt{d_{k}}})V_{i}W_{o,i+1}+\mathbf{x}_{i+1}$
Let $A_{i}=\frac{\mathbf{x}_{i}W_{q,i+1}K^{T}_{i}}{\sqrt{d_{k}}}$, $\hat{\delta}_{i}=\frac{\delta_{i}W_{q,i+1}K^{T}_{i}}{\sqrt{d_{k}}}$. By expanding Softmax to the first order:
$\mathbf{x}^{\prime}_{i+1}$ $=\big[\text{softmax}(A_{i})+J(A_{i})\hat{\delta}_{i}\big]V_{i}W_{o,i+1}$ (26)
$+\mathbf{x}_{i+1}+o(\hat{\delta}_{i}^{2}),$
where $J(A_{i})$ is the Jacobian of Softmax at $A_{i}$. We ignore the $o(\hat{\delta}_{i}^{2})$ remainder. (Xiao et al., 2024) has discovered that the pre-trained LLMs generally exhibit an “attention sink” feature, in which self-attentions heavily attend to the “sink token” at the beginning of a sequence. We adopt such feature to estimate $J(A_{i})||$. According to the statistical pattern concluded by (Xiao et al., 2024), attention logits $\mathbf{a}\in\mathbb{R}^{l}$ of a sequence with length $l$ is approximately:
$a_{i}=1\;(0<i<d),\;a_{j}=-1\;(i\geq d),\;d\ll l$ (27)
(Xiao et al., 2024) empirically adopts $d=4$. For a sequence of $l=1024$, $J(\text{softmax}(\mathbf{a}))||\approx 0.03\ll 1$, and decreases as $l$ increases. Therefore, we can conclude that the coefficient of the $\hat{\delta}$ term is significantly smaller than that of the Softmax term, making the contribution of the “depth factor” negligible. ∎
Appendix H Proof for Proposition 3.4
Proof.
In the forward pass of UniAttn:
$\mathbf{x}_{i+1}^{\prime}$ $=s_{i}(\mathbf{x}_{i+1}W_{v,i+1})W_{o,i+1}+\mathbf{x}_{i+1}$ (28)
$=s_{i}(\mathbf{x}_{i}+\delta_{i})W_{v,i+1}W_{o,i+1}+\mathbf{x}_{i+1}$
[…]
Appendix I More Insights on Linear Compensation